• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Jin Huang

    Log entry time 00:09:29 on November 07, 2009

    Entry number 297953

    keyword=Shift Summary Nov 06 Swing

    Shift start with control access. LHRS is ready. Vince and Ramesh are working in RHRS. RHRS diple is ramping up. RHRS momentum setting is 2.36GeV. VDC off on both arm. Empty target.

    16:22 Reset RHRS momentum to 2.63GeV/c; Cycling Q2/3

    16:50 MCC called that they are ready to tune beam through Hall A. Asking them to tune the beam to nominal beam position on white board: BPMA = (-2.0, 1.25)mm BPMB = (-3.0, 1.5)mm

    17:30 MCC can not get beam to requested position due to compton is now in beam. What we get is BPMA = (0.3,1.3) BPMB = (1.4,1.5). Continue to Carbon with hole target.

    17:33 Start detector check out run: 25269/4814

    17:40 Carbon hole data tell us we are far off on beam X position as we expected. Detector check out on both arm is on going. 

    17:48 Continue to moller measurement

    22:04 moller measurement done: Halog 297889

    22:14 BPM ->  BPMA = (-2.0, 1.25)mm BPMB = (-3.0, 1.5)mm after MCC applied some magic .....

    22:17 Target -> Carbon hole to confirm beam position. Run: 25271/4816 ........

    22:41 Established our golden beam position BPMA = BPMB = (-3.0, 2.0)mm, which perfectly center carbon hole and without beam angles.

    22:42 Harp scan again. There are Problems (Halog 297918). But we decide to move on to Q^2 measurements

    23:00 Q^2 Measurement : LD2 Target, Taking run at both L/R HRS. VDC on.   >200k Event each

    23:45 Try to deal with a BPM and high dead time problem on LHRS. See Halog 297942, 297946, 297948. Bob called in to help solve fix the DAQ.